OpenStudy (anonymous):

Find the zeros of the function. y=(x+4)(x-5) If possible tell me the steps too, not just the answer. Thanks! :) *Best answer gets medal*

OpenStudy (anonymous):

(x+4)(x-5)=0 this is true if x=-4 and x=5 so at values x=-4 and x=5 function is 0.

OpenStudy (anonymous):

Zeros are the values of x that make this function = 0 (x+4) = 0 f x=-4 (x-5) = 0 if x = 5. zero times anything is zero.

OpenStudy (anonymous):

I don't get it... My teacher's example showed foiling.

OpenStudy (anonymous):

"foiling"? These have to be the zeros: -4 and +5

OpenStudy (anonymous):

Okay, Thank you

OpenStudy (anonymous):

You are quite welcome.